OpenStudy (anonymous):

Find the slope of the cardioid r=2+2cos(theta) at the point corresponding to (theta)=pi/4

OpenStudy (irishboy123):

\(r=2+2cos(\theta)\) the slope takes you back into cartesian, ok? so you want \(\frac{dy}{dx}\) yep?

OpenStudy (anonymous):

This one specifically is nothing like the ones I've done in my previous assignment. So I'm confused as to how to even begin to tackle it.

OpenStudy (anonymous):

wouldn't the tangent of pi/4 in this case be the slope?

OpenStudy (irishboy123):

totally

OpenStudy (irishboy123):

do you know how to do that?

OpenStudy (irishboy123):

this is it https://www.desmos.com/calculator/qysgdfnvnr

OpenStudy (anonymous):

totally makes sense now that I have the graph! thanks!

OpenStudy (irishboy123):

cool!

OpenStudy (anonymous):

I have one more question. Maybe you can help with this one too, because MacLaurin Series are my weakest topic in this course so far.

OpenStudy (irishboy123):

just to complete this thread, the formula to get the slope in cartesian is \(\large \frac{dy}{dx} = \frac{\frac {d r}{d \theta} \ sin \theta + r \ cos \theta}{\frac {d r}{d \theta} \ cos \theta - r sin \theta}\) very deriveable....from the basic premises that \(x = r cos \theta \) etc

OpenStudy (irishboy123):

maclaurin no worries stick it in a new thread first, though